OpenStudy (anonymous):

Let R be the region in the xy-plane between the graph e^x and y=e^-x from x=0 and x=2. a. Find the volume of the solid generated when R is revolved about the x-axis.

OpenStudy (campbell_st):

 y = e^x y^2 = e^2x and e^-2x I'd use \[V = \pi \int\limits_{0}^{2} e^2x dx - \pi \int\limits_{0}^{2} e^{-2x} dx\]